OpenStudy (anonymous):

how do i intergrate -2 lxl +1

OpenStudy (anonymous):

\[ -2|x|+1 = \begin{cases} -2x+1&x>0\\ 2x+1&x<0 \end{cases} \]

OpenStudy (anonymous):

\[ \int _a^b-2|x|+1\;dx = \int_{a}^02x+1\;dx+\int_0^b-2x+1\;dx \]

OpenStudy (anonymous):

The antiderivative of \(|x|\) would be: \[ \begin{split} \int \begin{cases} x&x>0\\ -x&x<0 \end{cases} dx &= \begin{cases} \frac{x^2}{2}+C&x>0\\ -\frac{x^2}{2}+C&x<0 \end{cases}\\ &= C+\frac{x}{2}\cdot \begin{cases} x&x>0\\ -x&x<0 \end{cases}\\ &= \frac{x|x|}{2}+C \end{split} \]

OpenStudy (anonymous):

Using properties of integrals, we would say \[ \int -2|x|+1\;dx =\color{red}{ -2\int |x|\;dx}+\color{blue}{\int 1\;dx }= \color{red}{-x|x|} +\color{blue}{ x} +C \]

OpenStudy (anonymous):

if im calculating a definite integral from -1 to 1, I got my answer to be 2. would that be right?